Exchange rate is
0.98 USD per Canadian
150 Yen per Canadian
0.8 Euro per Canadian
20 Peso per Canadian

Watermelons are
6 US Dollars in Chicago
1200 Yen in Tokyo
7 Euros in Madrid
100 Pesos in Cancun

Which shows the most expensive juice?
Answer #1

Given, $ 0.98 = 1 Canadian Dollar

150 Yen = 1 Canadian dollar

=> 150 yen = 0.98 $

=> $ 1 = 150/0.98 = 153.0612 yen

0.8 Euro = 1 Canadian dollar

=> 0.8 € = 0.98 $

=> 1 $ = 0.8163 €

20 Peso = 1 Canadian dollar

=> 20 Peso = 0.98 $

=> 1 $ = 20/0.98 = 20.4081 Peso

In US watermelons cost $ 6

In Foto it costs 1200 yen

In $ terms it would be = 1,200/153.0612 = $ 7.84

In Spain it costs 7 euros

In $ terms it would be = 7/0.8163 = $ 8.175

In Mexico it costs 100 pesos. Convert it to $ we get

100/ 20.4081 = $ 4.9

7 euros in Madrid shows most expensive juice.

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me, please help. Thank you.